OpenStudy (anonymous):

Identify the cube root contained as a factor in each expression. I need someone to at least explain to me what the problem is asking me to do...I don't understand it (click to see problem)

OpenStudy (anonymous):

\[3\sqrt[3]{-54}, (3/2)\sqrt[3]{-250}, 4\sqrt[3]{-16}\]

OpenStudy (anonymous):

you have to find "nice" cube roots i can give you the first one: 3 * cube root of 2 * cube root of 27= 3 times cube root of 2 times 3 = 9 times cube root 2

OpenStudy (anonymous):

alright? ^^

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you're welcome